Shift Summary for DAY shift 8/30/99

Sami Amasha (samasha@hep.uchicago.edu)
Mon, 30 Aug 1999 15:30:16 -0500

Shift Summary for DAY shift, 30/Aug/99

On Shift: Sasha Ledovskoy, Sami Amasha

Hours of beam available at all intensities: 8.5
Hours of beam available at nominal intensity: 8.5
Hours of E832 data (special runs not included): 7.75
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: Changing tapes and initializing a run;
Laser illumination problems;
drt_fresh_start;

========================================================================
Runs ended this shift:
Run 14376 - 15.8M events - BEAM_832
Run 14377 - 6.0M events - BEAM_832 (Bad spills 76-end of Run: CsI quad
problems)

Tapes brought to FCC this shift:
(including tapes for online split)
QK*354-6,QKB123

========================================================================
Accesses during shift:
None.

========================================================================
Overview of shift:
Steady beam, continuous data taking.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14376 07:15 Inherit smooth E832 data taking.

07:45 Tapes full. Change tapes. Update ADGN constants.

14377 1 07:58 start run 14377: BEAM_832

08:22 notice while filling out run sheed that
vv'ewud has low efficiency

76 09:37 Pipeline errors-reset pipeline & do KFREND.

81 09:45 laser illumination problem: upper left quadrant
is not illuminated. Laser stability RMS and other
CsI plots have only 2325 entries. Paged Ed.
With the help of BobT, we decide to do a
drt_fresh_start.

09:57 End run 14377. Do drt_fresh_start.

14378 10:10 Start run 14378

11:39 During routine histogram check, we notice that
lslope vs. plane plot in dc#summary shows a frown:
1 point above red line, 4 others very close to line

11:56 SEM has been a little high recently. Call MCR.

141 13:18 ktevana errors. KFREND.

14:00 TRD 1-8 HV is ON

212 14:52 LIVE_TIME error. Stop triggers and check busy
status, restart triggers. Ktevana errors.
KFREND. Start triggers again during spill 214.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

o if we lose beam for >= 40 minutes, and the DA works, do a laser scan.

o keep your eyes on channel 450 (see log book p. 103). If it remains stable,
and we do a laser calibration, maybe we don't need to swap the divider
(ask Ed). If it is unstable, page Julie and Ed.